# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

### v2

@@ Abstract
Mathematical models are widely used to study ecological and epidemiological systems, but the accuracy of resulting inferences and forecasts depends on the estimation framework. This study compares Bayesian and Frequentist approaches across three biological models using four datasets: the Lotka–Volterra (LV) predator–prey system, a generalized logistic model (GLM) applied to lung injury and mpox data, and SEIUR epidemic model for COVID-19. To ensure a fair comparison, both approaches were implemented with a normal error structure. We first examined structural identifiability to determine which parameters can be recovered in principle. We then evaluated practical identifiability and fitting performance using four metrics: mean absolute error (MAE), mean squared error (MSE), prediction interval (PI) coverage, and weighted interval score (WIS). For the Lotka–Volterra model, we studied three types of observations: prey only, predator only, and both species together. The Frequentist workflow, implemented through QuantDiffForecast, uses nonlinear least squares and parametric bootstrap to quantify uncertainty. The Bayesian workflow, implemented through BayesianFitForecast, uses Hamiltonian Monte Carlo sampling to obtain posterior distributions and diagnostic measures. Our results show that Frequentist inference performs well when the system is fully observed data, as in the GLM or in the Lotka–Volterra model when both species are observed. Bayesian inference performs better when uncertainty is high and observations are limited or indirect, as seen in the SEIUR epidemic model. The identifiability analysis helps explain these differences by showing how observability shapes the reliability of parameter recovery and uncertainty quantification.
@@ 1 Introduction
The U.S. CDC’s FluSight Challenge used models to optimize influenza vaccine distribution and public health messaging [10, 11, 12]. During the West African and Democratic Republic of the Congo (DRC) Ebola outbreaks, models predicted transmission patterns and evaluated intervention effectiveness [13, 14, 15, 16, 17, 18]. More recently, forecasting models for mpox predicted its spread and assessed containment measures [19, 20, 21, 22].
@@
In ecology, predator–prey models help predict population dynamics and inform conservation strategies [23]. These applications demonstrate that reliable forecasts depend on accurate parameter estimation [24, 25, 26]. However, the ability to recover meaningful parameters from data critically depends on identifiability, that is, whether model parameters can be uniquely determined from available observations [27, 28, 29].
@@
Identifiability can fail when data are sparse, models are over-parameterized, or parameters are strongly correlated, leading to misleading inferences and unreliable forecasts. Recent work has shown that factors such as vaccination behavior, immunity duration, and data completeness shape both epidemic trajectories and the reliability of parameter estimates [30]. Understanding how identifiability interacts with estimation frameworks is essential to ensure that model-based forecasts are interpretable and reproducible. Structural or practical identifiability limitations can cause two inference frameworks to produce different results even under identical models and data.
@@
Two estimation paradigms dominate the field: Bayesian and Frequentist methods. Frequentist methods typically calibrate ODE models by optimizing a likelihood function or minimizing an objective function such as the sum of squared differences between observed and predicted values [31, 32, 33, 34]. These methods use algorithms such as gradient descent or the Levenberg–Marquardt algorithm, assume specific distributions for measurement errors (e.g., Gaussian, Poisson, or Normal), and quantify uncertainty through bootstrapping techniques [35, 36, 37, 38, 39, 40, 41, 42]. Frequentist methods are computationally efficient and often perform well when data are abundant and high quality [43, 25]. The QuantDiffForecast (QDF) toolbox implements this workflow for fitting models and generating predictions with quantified uncertainty [25].
@@
Bayesian methods, in contrast, apply Bayes’ theorem to combine prior distributions of parameters with the likelihood of observed data, producing posterior distributions that explicitly incorporate uncertainty [44, 45, 46, 47, 48, 49, 50, 51]. These methods typically use Markov Chain Monte Carlo (MCMC) algorithms to approximate posterior distributions, providing comprehensive measures of parameter uncertainty and prediction intervals [52, 53, 54, 55, 56, 57]. Bayesian methods can better navigate complex parameter spaces, avoid local optima, and handle incomplete or noisy data [58, 59, 60]. Tools such as Stan facilitate Bayesian estimation and forecasting, allowing rigorous uncertainty quantification and model validation [61, 62, 63, 64, 65].
@@
The BayesianFitForecast (BFF) toolbox implements this workflow with diagnostics such as the Gelman–Rubin $\hat{R}$ statistic [26]. Despite extensive use of both paradigms, prior comparisons often vary models, likelihoods, or preprocessing between methods, making it difficult to attribute observed differences to the estimation framework itself [66]. To address this gap, we conduct a controlled comparison of Bayesian and Frequentist inference under standardized conditions: the same models, the same normal error structure, and harmonized data preprocessing.
@@
We analyze three systems and four datasets representing different levels of complexity and observability: the Lotka–Volterra (LV) predator–prey model (Hudson Bay lynx–hare data), a generalized logistic model (GLM) for lung injury and the 2022 United States mpox outbreak, and an SEIUR epidemic model for the first COVID-19 wave in Spain [67, 68, 69]. For the Lotka–Volterra model, we analyze three observation scenarios (prey only, predator only, and both simultaneously) to assess how partial observability affects parameter recovery.
@@
We integrate structural identifiability analysis to determine which parameters can theoretically be recovered from the data, separating fundamental data limitations from algorithmic limitations [70, 71, 72]. We then evaluate practical identifiability and forecasting performance using four metrics: mean absolute error (MAE), mean squared error (MSE), 95% prediction interval (PI) coverage, and weighted interval score (WIS) [73, 74, 75, 76, 77].
@@
Our objectives are to (i) compare prediction accuracy across Bayesian and Frequentist inference, (ii) assess uncertainty calibration and diagnostics, and (iii) relate observed performance to structural versus practical identifiability under full versus partial observation. This provides practical guidance on when each paradigm is preferable based on data richness, observability, and uncertainty quantification needs [26].
@@
The remainder of this paper is organized as follows. Section 2 describes the four datasets analyzed in this study. Section 3 presents the mathematical models employed. Section 4 details the Bayesian and Frequentist inference methodologies. Section 5 presents the structural identifiability analysis. Section 6 reports the empirical results. The Discussion synthesizes our findings, and the Conclusion summarizes the main contributions and implications for biological modeling practice.
@@ 2 Data
We analyzed four datasets spanning ecological and epidemiological systems, each selected to represent different levels of data richness, temporal resolution, and observability (Table 1; Figure 1). All datasets consist of time-series observations at discrete time points $t_{n}$, where $n=1,\ldots,N$ denotes the observation index. The datasets include both population counts (for the ecological system) and incident case counts (for epidemic systems). Data preprocessing was standardized across all datasets to ensure comparability between inference methods; specific preprocessing steps are summarized at the end of this section.
@@ 2.1 Hudson Bay Lynx–Hare Data
The Hudson Bay lynx–hare dataset consists of annual counts of Canadian lynx (Lynx canadensis) and snowshoe hare (Lepus americanus) populations from 1900 to 1920, derived from Hudson’s Bay Company pelt return records [76, 77] . These records serve as a proxy for population abundance, with pelt counts reflecting relative population size. The dataset contains $N=21$ yearly observations for both prey (hare) and predator (lynx) populations. This dataset is particularly valuable for assessing parameter identifiability in predator–prey dynamics because it provides simultaneous observations of both interacting species over multiple population cycles. The data have been widely used as a benchmark for testing ecological models and parameter estimation methods. No smoothing, interpolation, or imputation was applied to the original data.
@@ 2.2 Lung Injury Data (EVALI)
This dataset comprises weekly incident case counts of electronic-cigarette or vaping product use-associated lung injury (EVALI) reported in the United States during 2019 [78, 79, 80] . The outbreak was first identified in mid-June 2019, with cases rapidly escalating through the summer and fall before declining by November 2019. We analyzed $N\approx 21$ weeks of data spanning mid-June to early November 2019, as reported through the Centers for Disease Control and Prevention (CDC) Morbidity and Mortality Weekly Report (MMWR) surveillance system. Case definitions followed CDC criteria for confirmed and probable EVALI cases. The data were aggregated by epidemiological week (MMWR week) with no backfill corrections applied beyond those included in the official CDC reports. This dataset represents a short-duration outbreak with a clear peak, making it suitable for testing phenomenological growth models under well-defined epidemic dynamics.
@@ 2.3 Mpox Data (United States, 2022)
We analyzed weekly incident confirmed or probable mpox cases in the United States during the 2022–2023 outbreak as reported in [81, 82]. The outbreak began in May 2022 and peaked in August 2022 before declining through early 2023. Data were compiled from CDC national surveillance reports aggregated by MMWR week. Case definitions followed CDC guidelines for confirmed (laboratory-confirmed orthopoxvirus with specimen typed as monkeypox virus) and probable cases (epidemiologically linked to a confirmed case without laboratory confirmation). The temporal window analyzed corresponds to the public reporting period documented in CDC surveillance dashboards (see references for access dates). This dataset provides an example of an emerging infectious disease outbreak with intensive surveillance and public health response, allowing evaluation of model performance during rapidly evolving epidemic conditions.
@@ 2.4 COVID-19 Data (Spain, First Wave)
This dataset consists of daily incident laboratory-confirmed COVID-19 cases reported during the first epidemic wave in Spain from February through May 2020 [83, 84] . Data were compiled from official reports by the Spanish Ministry of Health (Ministerio de Sanidad), with case dates based on report date rather than symptom onset date due to data availability constraints during the early pandemic response. The first wave in Spain exhibited rapid exponential growth followed by a peak in late March 2020 and subsequent decline following implementation of strict non-pharmaceutical interventions including a national lockdown. We restricted analysis to the first wave period to avoid complications from changing surveillance protocols, intervention policies, and emergence of new variants that occurred in later waves. Known reporting artifacts explicitly flagged in official data releases were excluded. This dataset represents a large-scale epidemic with significant underreporting and latent compartments (exposed and unreported infectious individuals), making it ideal for assessing inference methods under partial observability and model complexity.
Table 1: Summary of datasets analyzed in this study, including data source, temporal coverage, reporting frequency, and number of observations for each system.
Dataset Source & Period Frequency Observations (N)
@@
Lung Injury (EVALI, US) CDC MMWR reports (mid-Jun–Nov 2019) Weekly $\approx$21
@@
COVID-19 (Spain, 1st Wave) Spanish Ministry of Health (Feb–May 2020) Daily Variable
@@ 2.5 Preprocessing summary:
All datasets were processed using a standardized protocol to ensure comparability across inference methods. The preprocessing steps were as follows:
  1. 1.
    Temporal aggregation: The Hudson Bay lynx–hare data were used as published without aggregation. Lung injury and mpox data were aggregated at the weekly level (MMWR epidemiological week) as reported in official CDC surveillance. COVID-19 data for Spain were analyzed at the daily level as reported by the Spanish Ministry of Health.
  2. 2.
    Data type: The Hudson Bay dataset represents population abundance (pelt counts), while all epidemic datasets (lung injury, mpox, COVID-19) represent incident case counts per reporting period. For the generalized logistic model (GLM), cumulative case counts were derived by summing incident counts, and model predictions were compared to incident data via the time derivative of cumulative cases.
  3. 3.
@@
  5. 5.
    Temporal windows: Modeling windows for calibration and forecasting were selected based on epidemic phase and data availability. Specific windows for each dataset are described in the Results section.
  6. 6.
    Data quality control: For COVID-19 data, known reporting artifacts explicitly flagged in official Spanish Ministry of Health releases were excluded. For all other datasets, data were used as published in official surveillance reports without additional corrections.
@@ 3 Models
In this study, we employed three compartmental biological and epidemiological models to comprehensively assess the parameter identifiability of our two approaches: (i) the Lotka–Volterra model, describing predator-prey interactions and population feedback mechanisms; (ii) the Generalized Logistic Model, capturing flexible epidemic growth through nonlinear case dynamics; and (iii) the SEIUR model, explicitly tracking susceptible, exposed, infectious (reported and unreported), and recovered populations to account for both observed and hidden transmission. These models were selected to represent a progression in model complexity and data observability: from a fully observed ecological system (LV), to a parsimonious single-equation epidemic model (GLM), to a multi-compartment latent-state model with partial observability (SEIUR). This progression enables a systematic comparison of Bayesian and Frequentist inference performance across increasing levels of model structure and latent uncertainty.
3.0.0.1 Lotka-Volterra model
The Lotka–Volterra equations (Lotka 1925; Volterra 1926, 1927) consist of ODEs describing the population dynamics of two interacting species: one predator and one prey. Let $x$ denote the prey population and $y$ the predator population at time $t$. Volterra modeled the temporal dynamics of these populations as follows:
$\frac{dx}{dt}=\alpha x-\beta xy,\quad\frac{dy}{dt}=-\gamma y+\delta xy,$ (1)
with initial conditions
$x(0)=x_{0},\quad y(0)=y_{0},$ (2)
where $\alpha$ is the intrinsic growth rate of the prey population, $\beta$ is the predation rate coefficient, $\gamma$ is the mortality rate of the predator population, and $\delta$ is the predator’s growth rate per prey consumed. The initial conditions $x_{0}$ and $y_{0}$ specify the starting population sizes of prey and predator, respectively. The Lotka–Volterra model serves as a classical benchmark in population ecology, offering a well-understood test case for examining parameter identifiability and uncertainty propagation under full versus partial observability.
@@ 3.0.0.1 Lotka-Volterra model
Figure 2: Lotka–Volterra predator–prey diagram. circles represent the prey $x$ and predator $y$ populations. The self-loop on $x$ indicates intrinsic growth at rate $\alpha$. Curved arrows between $x$ and $y$ represent interactions: prey loss due to predation and predator growth from consuming prey. The downward arrow on $y$ represents natural mortality.
3.0.0.2 Generalized Logistic Model
The Generalized Logistic Model (GLM) is a flexible extension of the logistic function used to model S-shaped epidemic growth curves. It is given by
$\frac{dC}{dt}=r\,C^{p}(t)\left(1-\frac{C(t)}{K}\right),$ (3)
with initial condition
$C(0)=C_{0},$ (4)
where $C(t)$ is the cumulative number of cases at time $t$, $r$ is the generalized growth rate, $K$ is the final epidemic size, and $p\in[0,1]$ controls the growth dynamics: $p=0$ corresponds to constant incidence, $0<p<1$ to sub-exponential growth, and $p=1$ to exponential growth. Here, $C_{0}$ represents the first reported case. Moreover, $\frac{dC}{dt}$ serves as the observation operator mapping the model states to the reported incidence data. The GLM was chosen because it provides a parsimonious, single-equation description of epidemic trajectories and serves as a tractable case for testing inference accuracy and uncertainty when the full time series is directly observed. This model has also been widely applied in epidemiological settings to capture early and full-phase epidemic dynamics [85] . The GLM has been widely used as a phenomenological growth model to capture epidemic curves with rapid increase and saturation [85, 86] .
@@ 3.0.0.3 SEIUR Model
$\begin{gathered}\frac{dS}{dt}=-\beta_{f}(t)\frac{(I+U)S}{N},\quad\frac{dE}{dt}=\beta_{f}(t)\frac{(I+U)S}{N}-\kappa E,\quad\frac{dI}{dt}=\kappa\rho E-\gamma I,\\ \frac{dU}{dt}=\kappa(1-\rho)E-\gamma U,\quad\frac{dR}{dt}=\gamma(I+U),\quad\frac{dC}{dt}=\kappa\rho E,\end{gathered}$ (5)
with initial conditions
$(S(0),E(0),I(0),U(0),R(0),C(0))=(N-C_{0},0,C_{0},0,0,C_{0}),$ (6)
where $N$ is the total population, $\beta_{f}(t)$ is the time-dependent transmission rate, $\kappa$ is the incubation rate, $\gamma$ is the recovery rate, $\rho$ is the reporting proportion, and $C_{0}$ is the first reported case. The initial condition assumes that all initially infected individuals are reported, and the remaining population is susceptible. The time-dependent transmission rate $\beta_{f}(t)$ is defined as the following
@@
The time-dependent transmission rate $\beta_{f}(t)$ is modeled in a piecewise manner to reflect changes in population behavior and control measures following an intervention at time $t_{\mathrm{int}}$. Before the intervention, transmission is assumed to be constant at level $\beta_{0}$, representing relatively uncontrolled epidemic spread. After $t_{\mathrm{int}}$, transmission decreases smoothly toward a lower asymptotic level $\beta_{1}$ through an exponential decay, capturing the gradual adoption and increasing effectiveness of interventions such as mobility reduction, masking, or behavioral adaptation. The decay rate $q_{1}$ controls how rapidly these effects take hold, allowing the model to represent both abrupt and slow transitions while remaining parsimonious and identifiable.
This model represents a more realistic epidemic process with both observed and hidden states, making it ideal for testing how Bayesian and Frequentist methods handle latent variables, parameter coupling, and partial identifiability.
@@
Figure 4: Compartmental diagram of the SEIUR model with underreporting. Circles represent the epidemiological compartments. Solid arrows indicate transitions between compartments, and the dashed arrow indicates the source of observed cases.

### v3

@@ Abstract
Mathematical models are widely used to study ecological and epidemiological systems, but the accuracy of the resulting inferences and forecasts depends on the estimation framework. This study compares Bayesian and frequentist approaches across three biological models using four datasets: The Lotka–Volterra predator–prey system, a generalized logistic model (GLM) applied to lung injury and mpox data, and the susceptible-exposed-infected (reported and unreported)-recovered (SEIUR) epidemic model for COVID-19. To ensure a fair comparison, both approaches were implemented with a normal error structure. We first examined the structural identifiability to determine which parameters can be recovered in principle. We then evaluated practical identifiability and fitting performance using four metrics: Mean absolute error, mean squared error, prediction interval coverage, and the weighted interval score. For the Lotka–Volterra model, we studied three types of observations: Prey only, predator only, and both species together. The frequentist workflow, implemented through QuantDiffForecast, uses nonlinear least squares and parametric bootstraping to quantify uncertainty. The Bayesian workflow, implemented through BayesianFitForecast, uses Hamiltonian Monte Carlo sampling to obtain posterior distributions and diagnostic measures. Our results show that frequentist inference performs well when the system comprises fully observed data, as in the GLM or in the Lotka–Volterra model when both species are observed. Bayesian inference performs better when uncertainty is high and observations are limited or indirect, as seen in the SEIUR epidemic model. The identifiability analysis helps explain these differences by showing how observability shapes the reliability of parameter recovery and uncertainty quantification.
keywords:
Bayesian inference; frequentist inference; performance metrics; forecasting; identifiability
@@ 1 Introduction
The US center for Disease Control and Prrevention (CDC’s) FluSight Challenge used models to optimize influenza vaccine’s distribution and public health messaging [10, 11, 12]. During the West African and Democratic Republic of the Congo (DRC) Ebola outbreaks, models predicted the transmission patterns and evaluated the intervention’s effectiveness [13, 14, 15, 16, 17, 18]. More recently, forecasting models for mpox predicted its spread and assessed containment measures [19, 20, 21, 22].
@@
In ecology, predator–prey models help predict population dynamics and inform conservation strategies [23]. These applications demonstrate that reliable forecasts depend on accurate parameter estimation [24, 25, 26]. However, the ability to recover meaningful parameters from data critically depends on identifiability, that is, whether the model’s parameters can be uniquely determined from the available observations [27, 28, 29].
@@
Identifiability can fail when the data are sparse, the models are over-parameterized, or the parameters are strongly correlated, leading to misleading inferences and unreliable forecasts. Recent work has shown that factors such as vaccination behavior, immunity, and data completeness shape both an epidemic’s trajectories and the reliability of parameter estimates [30]. Understanding how identifiability interacts with estimation frameworks is essential to ensure that model-based forecasts are interpretable and reproducible. Structural or practical identifiability limitations can cause two inference frameworks to produce different results even under identical models and data.
@@
Two estimation paradigms dominate the field: Bayesian and Frequentist methods. Frequentist methods typically calibrate ODE models by optimizing a likelihood function or minimizing an objective function such as the sum of squared differences between the observed and predicted values [31, 32, 33, 34]. These methods use algorithms such as gradient descent or the Levenberg–Marquardt algorithm, assume specific distributions for measurement errors (e.g., Gaussian, Poisson, or normal), and quantify uncertainty through bootstrapping techniques [35, 36, 37, 38, 39, 40, 41, 42]. Frequentist methods are computationally efficient and often perform well when the data are abundant and of high quality [43, 25]. The QuantDiffForecast (QDF) toolbox implements this workflow for fitting models and generating predictions with quantified uncertainty [25].
@@
Bayesian methods, in contrast, apply Bayes’ theorem to combine the prior distributions of parameters with the likelihood of the observed data, producing posterior distributions that explicitly incorporate uncertainty [44, 45, 46, 47, 48, 49, 50, 51]. These methods typically use Markov chain Monte Carlo (MCMC) algorithms to approximate posterior distributions, providing comprehensive measures of parameter uncertainty and PIs [52, 53, 54, 48, 55, 56]. Bayesian methods can better navigate complex parameter spaces, avoid local optima, and handle incomplete or noisy data [57, 58, 59]. Tools such as Stan facilitate Bayesian estimation and forecasting, allowing rigorous uncertainty quantification and model validation [60, 61, 62, 63, 64].
@@
The BayesianFitForecast (BFF) toolbox implements this workflow with diagnostics such as the Gelman–Rubin $\hat{R}$ statistic [26]. Despite the extensive use of both paradigms, prior comparisons often vary models, likelihoods, or preprocessing between methods, making it difficult to attribute observed differences to the estimation framework itself [65]. To address this gap, we conduct a controlled comparison of Bayesian and frequentist inference under standardized conditions: The same models, the same normal error structure, and harmonized data preprocessing. We purposely use simple, low-dimensional standard models so we can make a clear, fair comparison and use differential algebra tools for structural identifiability.
@@
We analyze three systems and four datasets representing different levels of complexity and observability: The LV predator–prey model (Hudson Bay lynx–hare data), a generalized logistic model (GLM) for lung injury and the 2022 United States mpox outbreak, and an susceptible-exposed-infected (reported and unreported)-recovered (SEIUR) epidemic model for the first COVID-19 wave in Spain [66, 67, 68]. For the LV model, we analyze three observation scenarios (prey only, predator only, and both simultaneously) to assess how partial observability affects parameter recovery.
@@
We integrate a structural identifiability analysis to determine which parameters can theoretically be recovered from the data, separating fundamental data limitations from algorithmic limitations [69, 70, 71]. We then evaluate practical identifiability and forecasting performance using four metrics: Mean absolute error (MAE), mean squared error (MSE), 95% coverage, and the weighted interval score (WIS) [72, 73, 74, 75, 76].
@@
Our objectives are to (i) compare prediction accuracy across Bayesian and frequentist inference, (ii) assess uncertainty calibration and diagnostics, and (iii) relate observed performance to structural versus practical identifiability under full versus partial observation. This provides practical guidance on when each paradigm is preferable, depending on data richness, observability, and the need of uncertainty quantification [26].
@@
The remainder of this paper is organized as follows. Section 2 describes the four datasets analyzed in this study. Section 3 presents the mathematical models used. Section 4 details the Bayesian and frequentist inference methodologies. Section 5 presents the structural identifiability analysis. Section 6 reports the empirical results. The Discussion synthesizes our findings, and the Conclusion summarizes the main contributions and implications for biological modeling practice.
@@ 2 Data
We analyzed four datasets spanning ecological and epidemiological systems, each selected to represent different levels of data richness, temporal resolution, and observability (Table 1; Figure 1). All datasets consist of time-series observations at discrete time points $t_{n}$, where $n=1,\ldots,N$ denotes the observation index. The datasets include both population counts (for the ecological system) and incident case counts (for epidemic systems). Data preprocessing was standardized across all datasets to ensure comparability between inference methods; the specific preprocessing steps are summarized at the end of this section.
@@ 2.1 Hudson Bay lynx–hare data
The Hudson Bay lynx–hare dataset consists of annual counts of Canadian Lynx (lynx canadensis) and snowshoe hare (Lepus americanus) populations from 1900 to 1920, derived from Hudson’s Bay Company pelt return records [75, 76] . These records serve as a proxy for population abundance, with pelt counts reflecting the relative population size. The dataset contains $N=21$ yearly observations for both prey (hare) and predator (lynx) populations. This dataset is particularly valuable for assessing parameter identifiability in predator–prey dynamics because it provides simultaneous observations of both interacting species over multiple population cycles. The data have been widely used as a benchmark for testing ecological models and parameter estimation methods. No smoothing, interpolation, or imputation was applied to the original data.
@@ 2.2 Lung injury data (EVALI)
This dataset comprises weekly incident case counts of electronic cigarette or vaping product use-associated lung injury (EVALI) reported in the United States during 2019 [77, 78] . The outbreak was first identified in mid-June 2019, with cases rapidly escalating through the summer and fall before declining by November 2019. We analyzed $N\approx 21$ weeks of data spanning mid-June to early November 2019, as reported through the Centers for CDC Morbidity and Mortality Weekly Report (MMWR) surveillance system. Case definitions followed CDC criteria for confirmed and probable EVALI cases. The data were aggregated by epidemiological week (MMWR week) with no backfill corrections applied beyond those included in the official CDC reports. This dataset represents a short-duration outbreak with a clear peak, making it suitable for testing phenomenological growth models under well-defined epidemic dynamics.
@@ 2.3 Mpox data (United States, 2022)
We analyzed weekly incident confirmed or probable mpox cases in the United States during the 2022–2023 outbreak as reported in [79]. The outbreak began in May 2022 and peaked in August 2022 before declining through early 2023. Data were compiled from CDC national surveillance reports aggregated by MMWR week. Case definitions followed CDC guidelines for confirmed (laboratory-confirmed orthopoxvirus with the specimen typed as monkeypox virus) and probable cases (epidemiologically linked to a confirmed case without laboratory confirmation). The temporal window analyzed corresponds to the public reporting period documented in CDC surveillance dashboards (see the references for access dates). This dataset provides an example of an emerging infectious disease outbreak with intensive surveillance and public health response, allowing evaluation of the model’s performance during rapidly evolving epidemic conditions.
@@ 2.4 COVID-19 data (Spain, first wave)
This dataset consists of daily incident laboratory-confirmed COVID-19 cases reported during the first epidemic wave in Spain from February through to May 2020. Data were compiled from official reports by the Spanish Ministry of Health (Ministerio de Sanidad), with case dates based on the report date rather than the symptom onset date due to data availability constraints during the early pandemic response. The first wave in Spain exhibited rapid exponential growth followed by a peak in late March 2020 and a subsequent decline following the implementation of strict non-pharmaceutical interventions, including a national lockdown. We restricted analysis to the first wave period to avoid complications from changing surveillance protocols, intervention policies, and emergence of new variants that occurred in later waves. Known reporting artifacts explicitly flagged in official data releases were excluded. This dataset represents a large-scale epidemic with significant under-reporting and latent compartments (exposed and unreported infectious individuals), making it ideal for assessing inference methods under partial observability and model complexity.
Table 1: Summary of datasets analyzed in this study, including data sources, temporal coverage, reporting frequency, and number of observations for each system.
Dataset Source and Period Frequency Observations (N)
@@
Lung injury (EVALI, US) CDC MMWR reports (mid-June–November 2019) Weekly $\approx$21
@@
COVID-19 (Spain, 1first wave) Spanish Ministry of Health (February-May 2020) Daily Variable
@@ 2.5 Preprocessing summary:
All datasets were processed using a standardized protocol to ensure comparability across inference methods. The preprocessing steps were as follows.
  1. 1)
    Temporal aggregation: The Hudson Bay lynx–hare data were used as published without aggregation. Lung injury and mpox data were aggregated at the weekly level (MMWR epidemiological week) as reported in the official CDC surveillance. COVID-19 data for Spain were analyzed at the daily level as reported by the Spanish Ministry of Health.
  2. 2)
    Data type: The Hudson Bay dataset represents population abundance (pelt counts), while all epidemic datasets (lung injury, mpox, COVID-19) represent incident case counts per reporting period. For the GLM, cumulative case counts were derived by summing incident counts, and model predictions were compared with incident data via the time derivative of cumulative cases.
  3. 3)
@@
  5. 5)
    Temporal windows: Modeling windows for calibration and forecasting were selected on the basis of epidemic phase and data availability. Specific windows for each dataset are described in the Results section.
  6. 6)
    Data quality control: For COVID-19 data, known reporting artifacts explicitly flagged in official Spanish Ministry of Health releases were excluded. For all other datasets, the data were used as published in official surveillance reports without additional corrections.
@@ 3 Models
In this study, we used three compartmental biological and epidemiological models to comprehensively assess the parameter identifiability of our two approaches: (i) the LV model, describing predator-prey interactions and population feedback mechanisms; (ii) the GLM, capturing flexible epidemic growth through nonlinear case dynamics; and (iii) the SEIUR model, explicitly tracking susceptible, exposed, infectious (reported and unreported), and recovered populations to account for both observed and hidden transmission. These models were selected to represent a progression in model complexity and data observability: from a fully observed ecological system LV, to a parsimonious single-equation epidemic model GLM, and to a multi-compartment latent-state model with partial observability (SEIUR). This progression enables a systematic comparison of how Bayesian and frequentist inference across increasing levels of model structure and latent uncertainty. We chose these standard, low-dimensional models on purpose so the comparison is clear and fair and so the structural identifiability can be analyzed directly.
Lotka-Volterra model.
The LV equations (Lotka 1925; Volterra 1926, 1927) consist of ODEs describing the population dynamics of two interacting species: One predator and one prey. Let $x$ denote the prey population and $y$ the predator population at time $t$. Volterra modeled the temporal dynamics of these populations as follows:
$\frac{dx}{dt}=\alpha x-\beta xy,\quad\frac{dy}{dt}=-\gamma y+\delta xy,$ (3.1)
with the initial conditions
$x(0)=x_{0},\quad y(0)=y_{0},$ (3.2)
where $\alpha$ is the intrinsic growth rate of the prey population, $\beta$ is the predation rate coefficient, $\gamma$ is the mortality rate of the predator population, and $\delta$ is the predator’s growth rate per prey consumed. The initial conditions $x_{0}$ and $y_{0}$ specify the starting population sizes of the prey and predator, respectively. The LV model serves as a classical benchmark in population ecology, offering a well-understood test case for examining parameter identifiability and uncertainty propagation under full versus partial observability. Figure 2 illustrates the interaction structure of the LV predator–prey model.
@@
Figure 2: LV predator–prey diagram. Circles represent the prey $x$ and predator $y$ populations. The self-loop on $x$ indicates intrinsic growth at a rate $\alpha$. Curved arrows between $x$ and $y$ represent interactions: Prey loss due to predation and predator growth from consuming prey. The downward arrow on $y$ represents natural mortality.
Generalized logistic model.
The GLM is a flexible extension of the logistic function used to model S-shaped epidemic growth curves. It is given by
$\frac{dC}{dt}=r\,C^{p}(t)\left(1-\frac{C(t)}{K}\right),$ (3.3)
with the initial condition
$C(0)=C_{0},$ (3.4)
where $C(t)$ is the cumulative number of cases at time $t$, $r$ is the generalized growth rate, $K$ is the final epidemic size, and $p\in[0,1]$ controls the growth dynamics: $p=0$ corresponds to constant incidence, $0<p<1$ to subexponential growth, and $p=1$ to exponential growth. Here, $C_{0}$ is the number of reported cases on day 0. Moreover, $\frac{dC}{dt}$ serves as the observation operator mapping the model states to the reported incidence data. The GLM was chosen because it provides a parsimonious, single-equation description of epidemic trajectories and serves as a tractable case for testing the accuracy and uncertainty of inferences when the full time series is directly observed. This model has also been widely applied in epidemiological settings to capture early and full-phase epidemic dynamics [80] . The GLM has been widely used as a phenomenological growth model to capture epidemic curves with a rapid increase and saturation [80, 81]. Figure 3 illustrates the GLM and its observation process.
@@
$\begin{gathered}\frac{dS}{dt}=-\beta_{f}(t)\frac{(I+U)S}{N},\quad\frac{dE}{dt}=\beta_{f}(t)\frac{(I+U)S}{N}-\kappa E,\quad\frac{dI}{dt}=\kappa\rho E-\gamma I,\\ \frac{dU}{dt}=\kappa(1-\rho)E-\gamma U,\quad\frac{dR}{dt}=\gamma(I+U),\quad\frac{dC}{dt}=\kappa\rho E,\end{gathered}$ (3.5)
with the initial conditions
$(S(0),E(0),I(0),U(0),R(0),C(0))=(N-C_{0},0,C_{0},0,0,C_{0}),$ (3.6)
where $N$ is the total population, and $\beta_{f}(t)$ is the time-dependent transmission rate, $\kappa$ is the incubation rate, $\gamma$ is the recovery rate, $\rho$ is the reporting proportion. Here, $C_{0}$ is the number of reported cases on Day 0. The initial condition assumes that all initially infected individuals are reported and that the remaining population is susceptible. The time-dependent transmission rate $\beta_{f}(t)$ is defined as the following:
@@
The time-dependent transmission rate $\beta_{f}(t)$ is modeled in a piecewise manner to reflect changes in the population behavior and control measures following an intervention at time $t_{\mathrm{int}}$. Before the intervention, transmission is assumed to be constant at a level $\beta_{0}$, representing relatively uncontrolled epidemic spread. After $t_{\mathrm{int}}$, transmission decreases smoothly toward a lower asymptotic level $\beta_{1}$ through an exponential decay, capturing the gradual adoption and increasing effectiveness of interventions such as mobility reduction, masking, or behavioral adaptation. The decay rate $q_{1}$ controls how rapidly these effects take hold, allowing the model to represent both abrupt and slow transitions while remaining parsimonious and identifiable.
This model represents a more realistic epidemic process with both observed and hidden states, making it ideal for testing how Bayesian and frequentist methods handle latent variables, parameter coupling, and partial identifiability. Figure 4 shows the compartmental structure of the SEIUR model with under-reporting.
@@
Figure 4: Compartmental diagram of the SEIUR model with under-reporting. Circles represent the epidemiological compartments. Solid arrows indicate transitions between compartments, and the dashed arrow indicates the source of observed cases.